OpenStudy (anonymous):

If y varies directly as x and x= 8 when y=5, what is y when x = 9? Question 11 options: 72/5 45/8 40/9 6

OpenStudy (anonymous):

B

OpenStudy (anonymous):

45/8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

y = (5/8)x

OpenStudy (radar):

y=kx where k is the constant of proportionality. 5=k8 k=5/8 y=(5/8)9 y=45/8=5' 5/8